Give the equation of the line that passes through (F,G) and is perpendicular to the line that passes through (V,Z) (W,Z).
Answer by venugopalramana(3286)

Give the equation of the line that passes through (F,G)and is perpendicular to the line that passes through (V,Z) (W,Z).....FORMULA FOR SLOPE OF A LINE JOINING 2 POINTS (X1,Y1) AND (X2,Y2) IS (Y2-Y1)/(X2-X1).
HENCE SLOPE OF LINE JOINING (V,Z) AND (W,Z)IS..(Z-Z)/(W-V)=0
TWO LINES ARE PERPENDICULAR IF THE PRODUCT OF THEIR SLOPES(M1 AND M2)IS EQUAL TO -1...THAT IS M1*M2=-1
SO M*0=-1...OR M= INFINITY
A LINE WITH INFINITE SLOPE MEANS IT IS PERPENDICULAR TO X AXS OR PARALLEL TO Y AXIS.HENCE ITS EQN. IS OF THE FORM X=K..WHERE K IS A CONSTANT.
BUT THIS IS PASSING THROUGH (F,G)..HENCE (F,G) SHOULD SATISFY THIS EQN.
F=K
HENCE THE EQN.OF THE REQUIRED LINE IS Y=F